OpenStudy (anonymous):

Which factor do you cancel from the numerator and denominator to simplify x squared minus 4 times x minus 5 all over x squared minus 25.? a. x + 4 b. x − 4 c. x − 5 d. x + 5

OpenStudy (anonymous):

Start by factoring the denominator

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

\[\huge \frac{(x-4)(x-5)}{x^2-25}\] you can't cross anything out as of now so factor all the terms you can

OpenStudy (anonymous):

so the answer is b?

OpenStudy (anonymous):

Nope. You can keep guessing until you get a yeah or you can try and follow my instructions.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

a

OpenStudy (anonymous):

that isn't even a thing in the expression

OpenStudy (anonymous):

so, nah

OpenStudy (anonymous):

does this help? \[\huge \frac{(x-4)(x-5)}{x^2-25} = \huge \frac{(x-4)(x-5)}{x+5)(x-5)}\]

OpenStudy (anonymous):

yes its d!!!

OpenStudy (anonymous):

You can only cross out equivalent expressions. like so: